A ball of sticky gum launched indoors with an init
Answer #1

26) x component velocity = 15 * cos60 = 7.5 m/sec

27)   y component velocity = 15 * sin60 = 13 m/sec

28)   y component velocity =   sqrt(132 - 2 * 9.8 * 3)

                                             = 10.5 m/sec

29) amount of time =   (13 - 10.5)/9.8

                                  = 0.26 sec

30) horizontal displacement = 7.5 * 0.26

                                             = 1.9 m

31) Here, 1/2 * 5 * t2 = 20t

=>   t =   8 sec

=> far does car get = 20 * 8 = 160 m

32)   maximum distance = (10 * 10)/(1.6)

                                        =    63 m